Skip to main contentSkip to solution

What is the minimum number of times a fair coin must be tossed so that the probability of getting at least one head exceeds 0.95?

Solution

✅ Correct Option: 3

If the coin is tossed once, there will be 21=22^1 = 2 choices i.e. ={H,T}=\{H, \boxed{T}\}

If the coin is tossed twice, there will be 22=42^2 = 4 choices i.e. ={HH,HT,TH,TT}=\{HH, HT, TH, \boxed{TT}\}

If the coin is tossed thrice, there will be 23=82^3 = 8 choices i.e. ={HHH,HHT,HTH,HTT,THH,THT,TTH,TTT}=\{HHH, HHT, HTH, HTT, THH, THT, TTH, \boxed{TTT}\}

There will only be one case of "all tails" (as shown in the boxed outcomes above).


Using the complement:

P(at least one head)=1−P(all tails)P(\text{at least one head}) = 1 - P(\text{all tails})

The probability of "at least one head" must exceed 0.950.95, so:

1−P(all tails)>0.951 - P(\text{all tails}) > 0.95

P(all tails)<0.05P(\text{all tails}) < 0.05


Let nn = number of coin tosses

n=3n = 3: Probability of all tails =18=0.125=\dfrac{1}{8} = 0.125

n=4n = 4: Probability of all tails =116=0.0625=\dfrac{1}{16} = 0.0625

n=5n = 5: Probability of all tails =132≈0.03125=\dfrac{1}{32} \approx 0.03125

Since 0.03125<0.050.03125 < 0.05, the condition is satisfied.


Therefore, the minimum number of tosses is 55.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question